OpenStudy (lena772):

Choose the correct graph of the given system of equations. y + 2x = -1 3y - x = 4 graph of two lines that intersect at the point (negative 1, 1) graph of two lines that intersect at the point (1, 1) graph of two parallel lines with positive slopes None of the above

OpenStudy (shamil98):

1 + 2(-1) = -1 1 - 2 = -1 correct. 3(1) - (-1) = 4 3 + 1 = 4 So your answer is the first option.

hartnn (hartnn):

to find the point of intersection, you just need to solve those 2 equation simultaneously the values u get for x and y is your intersection point